OpenStudy (anonymous):

Simplify the complex faction!! Medal rewarded! 2/5g - 4/3g OVER 1/2g + 4/5g. Please explain!

OpenStudy (anonymous):

5g + 4 ( -5 +3g ) = 1 - g ; divide by 4 5g/4 -5 +3g = 1/4 - g/4 ; add like terms 17g/4 -5 = 1/4 - g/4 ; swap sides +g/4 +5 ---------------- 18g/4 = 21/4 ; and multiply by 4/18 g = 21/18